Under constant temperature, the volume occupied by a gas varies inversely , to the pressure applied. If the gas occupies a volume of 20 cubic inches under a pressure of 6 pounds per square inch, find the volume when the gas is subjected to a pressure of 10 pounds per squire inch.
Answer by Cromlix(4381)

Volume varies inversely as Pressure
V = k/P
(k = constant of variation)
Using first set of data to work out k
20 = k/6
k = 120
Formula: V = 120/P
V = 120/10
Volume = 12 pounds per sq inch.
Hope this helps.
:-)